OpenStudy (anonymous):

what is cot 0 in the triangle

OpenStudy (mathmate):

Hint: Use SohCahToa, and recall that cot(x)=\(\Large \frac{1}{tan(x)}\)

OpenStudy (anonymous):

so what is x

OpenStudy (mathmate):

x can be any value, but can apply to \(\theta\) in your diagram.

OpenStudy (anonymous):

oh ok ....1 min

OpenStudy (anonymous):

ok ....im confused as how to solve this but i have cot(3)=0.0524

OpenStudy (mathmate):

Do you know SohCahToa? Start with finding tan\(\theta\)!

OpenStudy (mathmate):

Neither tan nor cot equals 0.0524. You can find it directly from the diagram as a fraction, without using a calculator.

OpenStudy (anonymous):

so does tan=4/3

OpenStudy (anonymous):

tan=opp/adj right??

OpenStudy (mathmate):

Exactly! Now you can calculate cot(x)=1/tan(x)

OpenStudy (anonymous):

lol what is cot

OpenStudy (mathmate):

cot is cotangent, basically the reciprocal of tangent.

OpenStudy (anonymous):

so if it is the reciprocal then that means that cot is 3/4 ??

OpenStudy (mathmate):

Exactly!

OpenStudy (anonymous):

so 3/4(x)=4/3

OpenStudy (anonymous):

so the answer is 3/4

OpenStudy (mathmate):

That is correct!

OpenStudy (anonymous):

ok thanks for the help......your awesome

OpenStudy (mathmate):

You're welcome! :)